### v1

# From Semantic To Instance: A Semi-Self-Supervised Learning Approach

Abstract
Instance segmentation is essential for applications such as automated monitoring of plant health, growth, and yield. However, extensive effort is required to create large-scale datasets with pixel-level annotations of each object instance for developing instance segmentation models that restrict the use of deep learning in these areas. This challenge is more significant in images with densely packed, self-occluded objects, which are common in agriculture. To address this challenge, we propose a semi-self-supervised learning approach that requires minimal manual annotation to develop a high-performing instance segmentation model. We design GLMask, an image-mask representation for the model to focus on shape, texture, and pattern while minimizing its dependence on color features. We develop a pipeline to generate semantic segmentation and then transform it into instance-level segmentation. The proposed approach substantially outperforms the conventional instance segmentation models, establishing a state-of-the-art wheat head instance segmentation model with mAP@50 of $98.5\%$. Additionally, we assessed the proposed methodology on the general-purpose Microsoft COCO dataset, achieving significant performance improvement of over $12.6\%$ mAP@50. This highlights that the utility of our proposed approach extends beyond precision agriculture and applies to other domains, specifically those with similar data characteristics.
Keywords: Instance Segmentation, Semi-Supervised Learning, Self-Supervised Learning, Wheat Head Instance Segmentation, Precision Agriculture, Data Synthesis
1 Introduction
Semantic segmentation [32, 36] is the pixel-level classification task in which each pixel within an image is assigned to a predefined semantic category, such as sky, tree, or vehicle. In contrast, instance segmentation [14, 16] involves identifying and outlining individual objects with well-defined boundaries, even if they belong to the same semantic category.
Instance segmentation poses distinct challenges compared to semantic segmentation, as it demands precise identification and separation of individual objects, which is crucial for applications such as autonomous driving [41], medical imaging [8], and precision agriculture [11].
Instance segmentation offers a significant advantage in precision agriculture by providing precise identification of individual objects within an agricultural setting [3]. In contrast to semantic segmentation, it enables the differentiation of individual plants, fruits, or other agricultural product elements. This distinction is critical for tasks such as tracking and counting crops [10, 34], assessing plant health [42, 27], and monitoring growth patterns [42], resulting in more targeted interventions, better resource allocation, and improved yield predictions. However, training deep instance segmentation models is challenging due to the need for large-scale datasets with pixel-level annotations of each object instance and the significant computational demand involved.
Creating large-scale annotated datasets across various domains is often expensive, time-consuming, and requires rare and highly specialized expertise for data collection and annotation. Moreover, supervised learning introduces challenges such as annotation bias [4], lack of generalizability [40], and robustness [15]. Semi-supervised and self-supervised learning leverage partially labeled and unlabeled data, respectively, for model training. They present successful alternatives to supervised learning in addressing its inherent challenges. By reducing the reliance on manual data annotation, semi- and self-supervised learning approaches are greatly expanding the applicability of deep learning in domains where creating annotated datasets is impractical.
In precision agriculture [21], a data-driven approach to optimizing the monitoring and management of agricultural products, the application of semi- and self-supervised learning can improve the performance of deep learning techniques [26, 5, 35, 25]. However, the considerable variability in agricultural products and environments, influenced by factors such as weather conditions, lighting, and crop diversity, necessitates the creation of large and diverse annotated databases for effective model training, posing challenges even for semi- and self-supervised learning approaches [35].
In contrast to general-purpose datasets [33, 19, 16], agricultural data are often characterized by small, repetitive, and overlapping objects with hidden and fine boundaries [25, 6], making the manual pixel-accurate and object-specific annotations of such large-scale datasets infeasible.
This paper proposes a semi-self-supervised learning approach for instance segmentation of images featuring dense, self-similar, and overlapping objects resembling those found in precision agriculture. We experiment and evaluate the effectiveness of our proposed methodology on the wheat head instance segmentation task. In addition, we experimented with the well-known COCO dataset [18] to showcase the applicability of our proposed method in the general-purpose domains.
We first use a limited number of manually annotated images to generate large-scale synthetic datasets with pixel-accurate annotations for each instance. These datasets are then used to develop a instance segmentation model for segmenting images of wheat fields taken in uncontrolled environments. We also apply domain adaptation to address the domain shift between synthetic and real data. Moreover, we propose to use the input RGB image and its corresponding semantic segmentation mask to create an image-mask representation, referred to as GLMask, ultimately enabling the deep model to transform the semantic segmentation mask into an instance mask. The GLMask is constructed by combining multiple color spaces derived from the RGB format and a computationally generated segmentation map. More specifically, GLMask represents the Grayscale from the RGB channels, L channel of LAB, and a semantic segmentation Mask. Models that are overly reliant on color features are problematic in a diverse wheat dataset, where wheat plants change color due to growth stage and variable outdoor lighting. Therefore, the G-L components in  GLMask are designed to encourage the model to focus on shape, texture, and pattern while minimizing its dependence on color features.
Specifically, the proposed approach follows a semi-self-supervised learning paradigm, wherein only a small number of pixel-accurate and instance-specific human-annotated images and video frames are used to computationally generate two large-scale datasets. Then a high-performing YOLOv9 [39] model is trained in two consecutive pretraining and domain adaptation phases by leveraging the synthetic datasets. Additionally, we augment the supervisory signals, generated by a pretrained semantic segmentation model within the input data by incorporating GLMask as the model input. Furthermore, we experiment with pseudo-labeling, as a self-supervised approach, utilizing the labels generated by our pretrained model.
This work establishes a state-of-the-art top-view wheat head instance segmentation benchmark through the following contributions: (1) introducing a semi-self-supervised learning methodology for high-performance instance segmentation model development; (2) incorporating a shape/pattern emphasized image-mask representation for semantic to instance mask conversion, replacing conventional RGB images, which substantially enhances model performance by providing a strong self-supervisory signal; (3) creating a large-scale computationally annotated synthetic data for deep model training; and (4) a comparison of two domain adaptation methods, rotation augmentation and pseudo-labeling; (5) An evaluation of our proposed methodology on a general-purpose dataset for the instance segmentation task, demonstrating its applicability beyond domain-specific applications. Figure 1 provides an overview of our proposed semi-self-supervised learning approach. Nevertheless, the utility of our proposed approach extends beyond wheat crops and precision agriculture, making it applicable to any domain involving data with similar characteristics.
Figure 1: Overview of the proposed approach. A semantic segmentation model is used to generate the semantic mask of the image, along with the grayscale and L channel of LAB color maps, which are concatenated to generate the new image-mask representation to feed to the instance segmentation model.
2 Related Works
Precision agriculture employs advanced technology to conduct several agricultural analyses through both offline and online computer vision tasks [35]. In this domain, image segmentation is used to develop automatic tools [27, 37] and equip machines for monitoring and controlling crops, weeds, soil, as well as crop infection and growth in finer details [13, 12, 6, 24].
In this context, numerous studies have explored the utility of wheat head semantic segmentation in top-view images of real fields. Rawat et al. [30] experimented with the wheat head semantic segmentation task with a pool-based active learning approach, obtaining an Intersection over Union (IoU) of $0.70$ on only the 2020 version of the UTokyo subset of the Global Wheat Head Detection (GWHD) dataset [7]. Najafian et al. [25] proposed a semi-self-supervised approach for wheat head segmentation, generating synthetic data from limited annotations and employing domain adaptation, resulting in a Dice score of $0.91$. Myers et al. [24] enhanced synthetic data realism by training a CycleGAN to transform unrealistic synthetic data into more authentic images, achieving a Dice score of $0.796$ on the GWHD dataset [7] after training the model on the GAN-generated data and final applying pseudo-learning phase. In addition, Ghanbari et al. [12] studied the need for domain adaptation and data diversity creating large-scale simulated data in training deep models, evaluated specifically on a wheat head segmentation task. Ennadifi et al. [9] developed a DeepMAC [9] and YOLOv5 [38] model for single-head segmentation of wheat heads, individually extracted using wheat head bounding boxes from $701$ manually annotated images or YOLOv5 model-generated bounding boxes, achieved an average F1 score of $86\%$ with their best-performing model.
Instance Segmentation, a finer approach than semantic segmentation, goes beyond identifying regions and delineates individual object instances within each category [22, 11]. Particularly, recent advancements in foundation models for pixel-level classification tasks have greatly increased the use of instance segmentation across various domains, including precision agriculture, through fine-tuning or zero-shot learning. Particularly, given the challenges associated with data annotation, supervised fine-tuning of these pretrained models offers a more straightforward approach, which requires less data annotation, compared to training a model from scratch, which demands a substantial number of annotated samples. Furthermore, the availability of manually annotated datasets, such as GWHD [7], serves as another crucial component in optimizing pretrained deep learning models for the agricultural domain tasks.
Li et al. [17] utilized transfer learning to fine-tune SAM model [16] for crop disease and pest segmentation in agricultural domains. They employed three datasets: two for coffee-leaf diseases, incorporating $1,100$ images, and one for pest image segmentation, containing $5,464$ images of $10$ common agricultural pests. The coffee-leaf disease datasets were split into training and testing subsets with an $80/20$ ratio. The pest dataset was divided into two groups, each with five pest categories. Their final model achieved an IoU of $59.58\%$ on the coffee-leaf disease datasets and $89.56\%$ on the pest datasets.
Nayak et al. [27] fine-tuned a YOLOACT [1] instance segmentation model to locate and segment wheat spike regions infected by Fusarium Head Blight (FHB). They employed a transfer learning approach, utilizing a dataset comprising $1,800$ annotated side-view images, which included $600$ images of infected wheat fields. This work focused exclusively on segmenting small, infected spike regions within the side-view images without segmenting all spikes.
Gao et al. [11] developed a DL-based model for severity assessment of wheat FHB in four healthy, mild, moderate, and severe categories. To do so, two BlendMask models were trained on $2,832$ training and $922$ validation sub-image samples for FHB-diseased single-spike segmentation and $524$ training and $166$ validation whole spike segmentation of side-view images. The samples were captured in sunny, cloudy, rainy, etc. environmental conditions and with backgrounds such as soil, other spikes, sky, and clouds. However, in the whole wheat field image segmentation, only sharp images in the foreground were manually annotated and, thereby, segmented by the model. The final models achieve $48.23\%$ IoU for spike instance segmentation and a $52.41\%$ IoU for FHB disease detection. The severity assessment was conducted by comparing the segmentation map of the infected regions with the segmentation map of the entire spike.
Despite the similarities to the aforementioned studies, our work stands apart in several key aspects. We focus on the instance segmentation of entire wheat head objects within large-scale, high-resolution, and top-view wheat-field images characterized by dense, self-similar, and overlapping objects. In contrast to most current research that relies on supervised learning methodologies and extensive annotations, our approach leverages semi-self-supervised learning and minimal data annotation. However, similar to other works, we utilize pretrained models [16, 39] to facilitate the development process, avoiding the need to train such large models from scratch, and leveraging the transfer of their learned knowledge.
With regard to the advancements in state-of-the-art (SOTA) instance segmentation models, such as the Segment Anything Model (SAM) series [16, 29], we opted to utilize YOLOv9 [39]—recognized as the SOTA in object detection and localization, known for its accuracy and efficiency in instance segmentation—in this work. SAM, which has outperformed models including FastSAM [43], was extensively trained on over $1.1$ billion segmentation masks across 11 million licensed images. FastSAM, on the other hand, leverages YOLOv8 [31] and a CLIP [29] prompting head, achieving 50×50$\times$50 × better speed but with reduced performance compared to SAM. SAM2 [29] introduced memory mechanisms for real-time image and video segmentation. However, YOLOv9 stands out by introducing the integration of Programmable Gradient Information (PGI) to address the information bottleneck problem and proposing the Generalized Efficient Layer Aggregation Network (GELAN) for optimal parameter utilization and computational efficiency. This architecture’s capacity to flexibly integrate various computational blocks while maintaining high performance makes YOLOv9 particularly adaptable to a wide range of applications, including instance segmentation, aligning better with our work’s objectives.
3 Material and Methods
3.1 Data
We manually annotated 10 image frames, each sized $1080\times 1920$, taken from two top-view video clips of wheat fields from Ghanbari et al. [12] for data synthesis (section 3.2). According to  [23], these images represent two domains corresponding to different growth stages of wheat. Eight images, denoted as $\textit{W}_{tr}$, were from a Wheat Mature and Harvest-ready growth stage, while the remaining two images, denoted as $\textit{W}_{va}$, were from a Heading-complete growth stage. Figure 2 illustrates sample images from these two domains. Images in $\textit{W}_{tr}$ and $\textit{W}_{va}$ were used for synthesizing a training set and a validation set, respectively. Each frame contained $106.7$ wheat head instances on average with the number ranging from $14$ to $262$.
Additionally, we utilized $29$ background video clips, without any wheat plant, of which we used $15$ videos to extract a set of $50,072$ frames, referred to as $\textit{B}_{tr}$, for the synthesizing training set. From the remaining $14$ video clip, we extracted $13,253$ image frames (referred to as $\textit{B}_{va}$) for the validation set data synthesis. Images in $\textit{B}_{tr}$ and $\textit{B}_{va}$ were randomly cropped to a size of $1024\times 1024$.
Figure 2: Examples of video frames from the Harvest-ready growth stage ($\textit{W}_{tr}$, left) and Heading-complete stage ($\textit{W}_{va}$, right) wheat field domains, along with their corresponding human-annotated instance-specific masks (bottom row).
[…]
For model evaluation, we manually annotated a set of $365$ images, hereafter referred to as $\textit{GHD}_{te}$. The average number of wheat head instances for these images was $52.3$, ranging from 00 to $102$. Also, a set of 100 images of size $1024\times 1024$, referred to as $\textit{LateStage}_{te}$, were used for further model evaluation. Each frame in $\textit{LateStage}_{te}$ has an average of $58.12$ instances, ranging from $5$ to $146$. Figure 3 presents a few manually annotated examples from our evaluation sets.
We utilized the unannotated images from the GWHD dataset [7], which lacked instance annotations. These images have been partitioned into three subsets of $\textit{Pseudo}_{tr}$, $\textit{Pseudo}_{va}$, and $\textit{GHD}_{qa}$, with $3657$, $1476$, and $1382$ images, respectively [7]. We excluded the 365 images in $\textit{GHD}_{te}$ from $\textit{Pseudo}_{tr}$. We selected two external unannotated datasets for qualitative assessment: (1) $180$ samples, comprising $10$ samples from each of the $18$ domains from $\textit{GHD}_{qa}$ and (2) $100$ video frames, with $10$ image frames extracted from $10$ different videos representing distinct wheat field domains, denoted as $\textit{HandHeld}_{qa}$. These videos were captured with high-resolution handheld cameras and were center-cropped to $1024\times 1024$ pixels.
Figure 3: Examples illustrating the diversity within our human-annotated test sets. The top row represents samples from our single-domain test set $\textit{LateStage}_{te}$. The bottom row shows samples from our $18$ domains test set $\textit{GHD}_{te}$.
3.2 Data Synthesis
We adopted the cut-and-paste methodology originally proposed by [26] for object detection and applied it to the instance segmentation data synthesis with some modifications. As the primary adjustment, we increased the number of manually annotated frames—containing larger wheat head objects—to $10$, categorized as $\textit{W}_{tr}$ and $\textit{W}_{va}$. Additionally, we incorporated a wider range of video frames from various background scenes, sourced from $\textit{B}_{tr}$ and $\textit{B}_{va}$, to enhance the final synthesized data variability.
[…]
Figure 4: Data synthesis pipeline for instance segmentation, adapted from [25], extracts fake and real wheat heads from manually-annotated images (detailed in [25]), categorized into Harvest-ready ($\textit{W}_{tr}$) and Heading-complete ($\textit{W}_{va}$) stages subsets. Background frames were similarly extracted, diversified, and categorized into $\textit{B}_{tr}$ and $\textit{B}_{va}$. Fake ($\textit{F}_{tr}$ and $\textit{F}_{va}$) and real ($\textit{R}_{tr}$ and $\textit{R}_{va}$) wheat heads were overlaid on the backgrounds, with smaller heads used when overlaying more than $50$ objects and larger heads used for fewer objects, synthesizing field images captured at various altitudes with the corresponding instance segmentation maps.
To synthesize each sample, we randomly selected $10$ to $100$ fake and then real wheat heads to overlay on the backgrounds. However, dissimilar to [25], we divided the real wheat heads into two subsets based on the object’s larger dimensions, with smaller and larger wheat heads in a $50\%-50\%$ ratio. Overlaying $10$ to $50$ objects, we selected from the subset containing larger heads; for more than $50$ overlying objects, we chose from the subset with smaller wheat heads. This approach aimed to enhance the realism of the synthesized samples by simulating wheat field images captured from varying altitudes. Additionally, we designed our pipeline to computationally generate the instance segmentation maps and contour labels, required for YOLO models, in parallel with synthesizing each image, for model development.
For training with the YOLOv9 [39] model, we chose not to apply additional online augmentations, relying solely on the augmentations integrated into YOLOv9. Instead, we increased the variability of our dataset by applying strong augmentations to the synthesized samples, in an offline manner. Specifically, we applied Flip, Rotation, and Elastic transformations for fake and real wheat head objects before overlying them. Furthermore, we opted for a long list of pixel-level augmentations including Color Jitter, Channel Shuffle and Dropout, Solarize, Blur, and Noise for augmenting the synthesized images at the end of the process. We employed the Albumentations [2] Python package, version 1.4.14, in implementing all the augmentation transformations.
3.3 Rotation-Augmented Data
For the second phase of the model training, we utilized real wheat field images to implement our domain adaptation method. As previously discussed, pixel-accurate instance annotation of many agricultural images is infeasible, and the $19$ samples in $\textit{Rotated}_{tr}$ were insufficient for fine-tuning a large model such as YOLOv9. Addressing these limitations, we extended the $\textit{Rotated}_{tr}$ dataset by applying image rotation. To enhance variability in our limited training dataset and better represent natural environments, we utilized rotation to simulate various angled top-down views of bending wheat stems, particularly effective for aerial imaging in windy conditions. Following the method used by [25], each image/mask pair was rotated with 00 to $259$ degrees, resulting in a large-scale dataset of $6840$ samples. To maintain the authenticity of samples for addressing domain shifts between synthetic and real data, we chose to exclude additional pixel-level or spatial-level transformations, apart from those applied by the YOLOv9 models in an online manner. We repeated the same process for the $\textit{Rotated}_{va}$ dataset, expanding it to match the sample size of $\textit{Rotated}_{tr}$.
Table 1: Summary of the datasets utilized, manually annotated, or computationally generated.
Dataset Aim Notation
Synthetic Train $\textit{SYN}_{tr}$
Valid $\textit{SYN}_{va}$
Rotation-Augmented Train $\textit{Rotated}_{tr}$
Valid $\textit{Rotated}_{va}$
Pseudo-Labeled Train $\textit{Pseudo}_{tr}$
Valid $\textit{Pseudo}_{va}$
Manually Annotated Test $\textit{LateStage}_{te}$
Test $\textit{GHD}_{te}$
Unannotated Quality Assessment $\textit{HandHeld}_{qa}$
Quality Assessment $\textit{GHD}_{qa}$
3.4 GLMask Generation
We utilized the Grayscale (G) and the perceptual lightness channel of LAB (L) color space versions of each RGB image, along with the corresponding semantic mask (M), to create a new image-mask representation of the RGB image, referred to as GLMask. We used pseudo-labels generated by the model developed for wheat head semantic segmentation by Najafian et al. [25]. By embedding the semantic mask, we aimed at efficient segmentation transfer from semantic to instance masks. Additionally, we designed GLMask to address a common issue in convolutional neural networks, where models often become overly reliant on color features. This reliance can hinder model generalization in the context of a diverse wheat dataset, as wheat color varies significantly due to growth stages and changing outdoor lighting conditions. By incorporating grayscale and the L channel of the LAB color spaces, GLMask intentionally reduces color information, encouraging the model to prioritize shape, texture, and pattern features.
The grayscale image $G$, derived by weighting the RGB images using $G=0.2125\times R+0.7154\times G+0.0721\times B$ formula [28], represents brightness, which is a subjective visual perception influenced by both luminance and color. It enhances contrast and improves edge and shape visibility. $G$ also reduces color-induced noise, thereby improving the quality of agricultural images, which generally contain higher noise levels than general-view images.
On the other hand, the L channel of the LAB color space represents an objective, measurable quantity of perceptual lightness. In contrast to grayscale, it approximates human vision perceptually more accurately in a uniform manner. The L channel separates luminance from chromatic information, improving shadow identification and highlighting shape and structural details. It provides a more balanced representation of light and dark areas while reducing the color-related artifacts. To create the three-channel GLMask image (please see Figure 1), we concatenated G, L, and M, where M contains values of 0 or 255. GLMask is used as the input for our model in all the training or evaluation phases.
3.5 Pseudo-Labeling
We explored pseudo-labeling as an alternative to rotation augmentation utilized for domain adaptation. This widely used self-supervised approach enhances model generalization by leveraging larger pseudo-labeled datasets, improving performance and reducing overfitting, while also lowering labeling costs, particularly in domains such as precision agriculture with packed objects. We utilized the best-performing model in the first phase, trained on the synthetic data, to label the unannotated $\textit{Pseudo}_{tr}$ and  $\textit{Pseudo}_{va}$ sets. Subsequently, we used these sets to fine-tune the model, initially trained on only synthetic data. We compare the performance of this model with the model developed using $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$ datasets to obtain our best-performing model as a better domain adaptation technique.
3.6 Experiments
As described in section 2, we selected the pretrained extended model YOLOv9-Seg architecture with $60,512,800$ parameters for all our experiments. We trained and evaluated the model using the data outlined in Table 1, with training conducted for $100$ epochs in each training phase on $1024\times 1024$ images. Using the AdamW optimizer [20], we set an initial learning rate of $31e-$ 3 to train the model on the $\textit{SYN}_{tr}$ and $\textit{SYN}_{va}$ datasets, resulting in SynModel, and a learning rate of $41e-$ 4 to fine-tune the $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$ datasets, leading to the development of RoAModel. We also applied a learning rate of $41e-$ 4 and used the $\textit{Pseudo}_{tr}$ and $\textit{Pseudo}_{va}$ datasets to conduct the pseudo-labeling experiment by fine-tuning model SynModel in comparison to model RoAModel, yielding PseModel. We further developed a baseline model, BaseModel, identical to SynModel but on RGB images. The batch size was set to $18$ and the IoU to $0.7$. The deterministic and seed parameters were set to ensure reproducibility, and the remaining parameters were kept at their default values. The default data augmentations applied included HSV color augmentation, horizontal flip, mosaic augmentation, erasing, translation, and scaling.
We subsequently trained two models on the Microsoft COCO 2017 dataset, using its training and validation sets, with the same exact parameters as those utilized in the wheat experiments with some exceptions. These models were trained on both the RGB and GLM versions of the dataset, with an adjusted batch size of $48$, enabled by a training image size of $640\times 640$ (up from the batch size of $18$ for $1024\times 1024$ images in wheat experiments). Given the dataset’s substantial size, both COCO models were trained for $50$ epochs. Additionally, we set IoU and conf parameters to be 0.6 and 0.25, respectively, during the evaluation. These two parameters enhanced both RGB and GLM model performances on the COCO validation set. We evaluate the models on the validation set of this dataset, as annotations for its test subset were unavailable. Note that we only require the binary (foreground vs. background) mask to create the GLM version of the data. This implies that, unlike the binary mask used as the semantic mask for the single-class wheat dataset, it serves as a binary mask instead of the semantic mask for the COCO dataset, which contains 80 distinct object classes.
All experiments were executed on three GPUs, specifically two A40 GPUs each equipped with $48$ GB GDDR6, and one Tesla V100S PCIe with $32$GB GPU. The environment was configured with Python 3.12, Ultralytics version 8.2.79, and PyTorch 2.4.0+cu124.
4 Results
Table 2: Performance evaluation of SynModel, RoAModel, and PseModel on our manually annotated test sets, $\textit{LateStage}_{te}$ and $\textit{GHD}_{te}$. We utilized YOLOv9e-Seg [39], denoted as YOLOv9e, as the base pretrained model for our experiments. The Color column indicates whether the experiments were conducted on the RGB version of the data or using our approach with GLMask data.
Model Pretrained Model Test Data Color Space P $\uparrow$ R $\uparrow$ mAP@50 $\uparrow$ mAP@50-95 $\uparrow$
BaseModel YOLOv9e $\textit{LateStage}_{te}$ RGB 81.9 38.2 49.7 24.9
$\textit{GHD}_{te}$ RGB 66.9 41.9 50.4 27.1
SynModel YOLOv9e $\textit{LateStage}_{te}$ GLM 86.5 82.0 88.2 61.2
$\textit{GHD}_{te}$ GLM 95.2 95.3 97.9 77.0
RoAModel SynModel $\textit{LateStage}_{te}$ GLM 93.2 93.3 95.7 80.7
$\textit{GHD}_{te}$ GLM 98.1 96.5 98.5 85.3
PseModel SynModel $\textit{LateStage}_{te}$ GLM 77.1 71.7 77.5 49.0
$\textit{GHD}_{te}$ GLM 95.8 95.2 98.0 76.6
Table 2 presents the performances of the developed models using Precision, Recall, and mAP metrics on our test sets. The baseline model BaseModel that was trained on the RGB version of the synthetic data ($\textit{SYN}_{tr}$ and $\textit{SYN}_{va}$) performed significantly lower than the SynModel, which was trained under the same condition but on the GLMask version of the synthetic data.
Figure 5: Prediction Performance of RoAModel across the $18$ domains of $\textit{GHD}_{te}$ test set.
SynModel, trained solely on synthetic data ($\textit{SYN}_{tr}$ and $\textit{SYN}_{va}$), obtained mAP@50 of an $88.2\%$, yielding $38.5\%$ improvement over BaseModel, on the $\textit{LateStage}_{te}$, representing a single data domain. Despite being trained only on synthetic data, the model achieved $47.5\%$ higher performance in comparison toBaseModel with $97.9$ mAP@50 on the $\textit{GHD}_{te}$, which is considered an external test set for this model. Nevertheless, the performance of both BaseModel and SynModel on the $\textit{GHD}_{te}$, which serves as an external evaluation set for both models, demonstrate the efficiency of our data synthesis approach for instance segmentation.
RoAModel was built upon SynModel by being fine-tuned on the rotation-augmented dataset $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$. RoAModel obtained the high-performance of $95.7\%$ mAP@50 on our single-domain $\textit{LateStage}_{te}$ set and a mAP@50 of $98.5\%$ on $\textit{GHD}_{te}$. Figure 5 displays the prediction performance of this model on samples from all $18$ domains of $\textit{GHD}_{te}$. Considering the better mAP@50-95 scores of this model on both sets, Figure 6 visually presents the prediction quality of the RoAModel in comparison to BaseModel on a few samples chosen from $\textit{LateStage}_{te}$ and $\textit{GHD}_{te}$ test sets. Regarding the quality assessment on $\textit{HandHeld}_{qa}$ and $\textit{GHD}_{qa}$oAModel visually demonstrates substantial accuracy across all $28$ data domains (please see Figures 7 and  8). We conducted an additional comparative experiment by training PseModel on SynModel, using $\textit{Pseudo}_{tr}$ and $\textit{Pseudo}_{va}$. This model underperformed RoAModel when evaluated on both test sets.
Figure 6: Qualitative performance comparison between the BaseModel and our best model RoAModel. The first column presents an image from $\textit{LateStage}_{te}$ while the rest of the columns display samples chosen from $\textit{GHD}_{te}$. The prediction instance masks are overlaid on the images.
Figure 7: Prediction performance of RoAModel on across $18$ domains within one of our external quality assessment set, $\textit{GHD}_{qa}$.
Figure 8: Prediction quality of RoAModel across the $10$ domains of our external handheld-captured quality assessment set, $\textit{HandHeld}_{qa}$.
Table 3: Performance comparison of the same YOLOv9e architectures trained from scratch using RGB (base model) and GLM (our approach) versions of COCO dataset [19].
Model Pretrained Model Color Space P $\uparrow$ R $\uparrow$ mAP@50 $\uparrow$ mAP@50-95 $\uparrow$
CocoModel None RGB 60.0 45.5 55.4 37.6
None GLM 70.6 58.7 68.0 55.4
We additionally experimented with our proposed approach on a general-purpose dataset and task using the Microsoft COCO 2017 dataset [19], extending beyond precision agriculture to asses it’s applicability across broader domains. We trained two models, one using the dataset’s RGB version of training and validation sets and one applying our proposed GLM version. Table 3 presents model performances on the validation set of this dataset, illustrating that our proposed approach achieved improvements in up to $12.6\%$ of mAP@50 and over $17.8\%$ in mAP@50-95. Additionally, Figure 9 illustrates a comparison of prediction performances between the RGB and GLM models, displayed in the second and third rows, respectively, with the ground truth shown in the first row. Each column in the figure represents a distinct sample, annotated with the ground truth labels, annotations generated by the base RGB model, and annotations produced by the GLM model developed using our proposed approach.
Figure 9: Prediction performance of the COCO models on the Microsoft COCO 2017 dataset [18]. Our proposed GLM approach consistently achieved superior segmentation performance (columns B, C, D, and E), and obtained higher confidence scores for detected objects (A through F). In some cases both RGB and GLM models failed to detect objects of interest, examples include the truck in column D and people in F. In addition, near-perfect segmentation performance was observed in cases such as A.
5 Discussion and Limitations
In this work, we proposed a semi-self-supervised learning approach for the instance segmentation task of images containing densely overlapped objects, specifically evaluating the methodology in agricultural domains with a focus on wheat head segmentation. Utilizing only $10+36$ manually annotated images, we developed a high-performing model that demonstrated strong performance across diverse wheat domains, each characterized by distinct crop phenotypes, canopy structures, and vegetation [7]. We selected the YOLOv9 [39] model due to its high capacity, speed, and efficiency in image segmentation, making it particularly well-suited for deployment on agricultural machinery performing real-time tasks. It is also important to note that the real-time YOLO architectures are widely recognized as state-of-the-art in instance segmentation, making them a robust and reliable baseline model for our approach.
We illustrated that human instance segmentation annotation of wheat field images and video frames is both challenging and time-consuming. To address this, we generated two large-scale datasets including synthetic, large-scale rotation-augmented, along with their computationally labeled instance masks, to bypass the laborious manual annotation process. Additionally, we introduced GLMask, a novel image representation technique combining grayscale, the L channel of the LAB color map, and computationally or model-generated semantic masks. GLMask replaces RGB inputs in the instance segmentation models development, enhancing semantic to instance segmentation transfer while reducing the model’s over-reliance on color information.
Using GLMask, SynModel, which was trained exclusively on synthetic data, achieved an impressive $97.9$ $mAP@50$ on the external test set of $18$ diverse wheat domains collected around the world, further demonstrating the effectiveness of the proposed data synthesis method in instance segmentation model training. Furthermore, it demonstrates the effectiveness of our proposed approach in enhancing the model’s generalizability, allowing it to adapt to various data domains despite being trained on only one domain. To enhance model robustness and generalizability to real data, we implemented a domain adaptation technique, resulting in almost $1\%$ of mAP@50 and over $8\%$ increase in mAP@50-95. We further explored pseudo-labeling as an alternative to rotation-augmented methodology. However, the results indicated that this approach was less effective in comparison to the rotation-augmented domain adaptation technique.
Beyond the agricultural domain, evaluating our proposed approach on a general-purpose instance segmentation task using the MS COCO dataset [19] supports its applicability across diverse fields. Although only binary masks, rather than semantic masks, were used to generate the GLM version of the COCO dataset for model development, this approach achieved a $17.8\%$ improvement in performance compared to the baseline YOLO model trained on the RGB version under identical training and evaluation conditions for fair comparison. As our approach originally advocated, the generation of GLM images using semantic masks instead of binary masks could further enhance CocoModel performance, suggesting that future studies should explore the use of semantic masks and additional color spaces to create GLM images.
Figure 10: (left) Large objects occluded by other objects resulting in the misidentification of fragmented parts as separate objects. (right) High-altitude images with numerous small wheat heads where overlapping objects are erroneously labeled as a single wheat head.
Despite the advancements, our model exhibits limitations in specific scenarios. For example, when large objects are occluded by another object and become fragmented, the model misidentifies the separated parts as distinct objects. Additionally, in images captured from high altitudes with numerous small wheat heads, the model may merge overlapping objects into a single label, leading to inaccuracies in object identification. These phenomena are visually demonstrated in Figure 10. We suggest further domain-specific model training to address these limitations.
In this study, we opted to use the YOLO model, a real-time state-of-the-art model for instance segmentation, to conduct our experiments. This choice aligns with our objective of developing precision agriculture models deployable on agricultural machinery, capable of real-time predictions on larger-scale images, such as $1024\times 1024$. However, this study lacks an evaluation of our proposed approach on other model architectures, such as SAM variations [16, 29], which, while resource-intensive for training and inference, may yet have limited applicability across various domains.
6 Conclusion
We proposed a semi-self-supervised learning approach for transforming semantic to instance-level segmentation, utilizing a semantic segmentation model to generate semantic masks. We then developed a pipeline that takes an image and its semantic mask as input and produces an instance-level segmentation. We showed the utility of our approach for wheat head instance segmentation; however, our approach could be utilized for similar tasks from other domains as well, as demonstrated through experiments on the COCO dataset.